1))))A person’s lungs might hold 6.0 x 10-3 m3 of air at body temperature (37 0C) and atmospheric pressure (101 kPa). Given that air is 21% oxygen (O2), find the number of O2 molecules in the lungs

ans: 2.97 x 1022

2))

If the person takes a particularly deep breath, so that the lungs hold a total of 1.5 x 1023 molecules, what is the new volume of the lungs?

[Answer: 6.35 x 10-3 m3 ]
Answer #1

1)V = 6*10^-3 m^3, T =37 oC =37+273 = 310 K, P =101 kPa

From ideal ga equation

PV = nRT

101*1000*6*10^-3 = n*8.314*310

n = 0.235 moles

only 21% oxygen

n = N/Na

N = 0.21*0.235*6.023*10^23

N = 2.97*10^22

(2) N = 1.5*10^23

PV = (N/Na)RT

101*1000*V = (1.5*10^23/(6.023*10^23))*8.314*310

V = 6.35*10^-3 m^3